f(x)=x^2-5x+3
part 1: f(x+h)
part 2: f(x+h)-f(x) over h ,h does not equal 0

Answer by RAY100(1637)

f(x) = (x)^2 -5(x) +3
.
f(x+h) = (x+h)^2 -5(x+h) + 3
.
f(x+h) = (x^2 +2xh +h^2) -5x-5h + 3
.
f(x+h) - f(x) = h^2 +2xh -5h
.